OpenStudy (anonymous):

Q and r are independent events. Find p(q and r) when p(q)=0.4 and p(r)=0.15

OpenStudy (anonymous):

P(q)=.4 P(r)=.15 P(Q)=4/5 , P(R)= 4/11

OpenStudy (anonymous):

Where did the p(q)=4/5 and the other come from? :/ I'm sorry, I don't understand!

OpenStudy (anonymous):

sorry i meant, P(Q)=4/10 , P(R)= 4/26.66666666

OpenStudy (anonymous):

Okay, can you please explain to me on how to get there? :/

OpenStudy (anonymous):

honestly i just kept dividing random numbers by 4 till i got .4 and .15, the first reply i looked it up but then after i did it I realized it was incorrect.

OpenStudy (anonymous):

Oh okay, okay :) thank you!!

OpenStudy (anonymous):

any time.

OpenStudy (anonymous):

any time.